TRIPLE TRANSITIONS IN SOLID HYDROGEN AND DEUTERIUM

\underline{M. MENGEL}, B. P. WINNEWISSER AND M. WINNEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, Heinrich-Buff-Ring 58, D-35392 Giessen, Germany.

The infrared absorption spectrum of the solid hydrogens is well understood in terms of a pairwise multipolar induction mechanism leading to the infrared activity of single and double transitions. However, in recent experiments we have identified various triple transitions for the first time ^{a} ^{b}. There are two basic mechanisms for the simultaneous excitation of three hydrogen molecules upon absorption of one photon of radiation:

• Intensity transfer due to mixing of states.

• Three-body induced dipole moments.

The infrared activity of triple transitions due to intensity transfer resulting from mixing of states was suggested by Tipping et al. ^{c} As examples of this mechanism we present the transitions Q_{1}(0) + Q_{1}(0) + S_{0}(0) and S_{1}(0) + Q_{1}(0) + S_{0}(0) in both solid hydrogen and deuterium. However, with this concept it is not possible to explain the triple transition S_{1}(0) + Q_{1}(0) + Q_{1}(0) which we have also observed in both isotopic species of the hydrogen crystal. This transition is located at 12 788 cm^{−1} in solid para-H_{2} and at 9123.5 cm^{−1} in solid ortho-D_{2}. We present a theory of three-body induced dipole moments in order to account for the linestrength of this very remarkable transition. Experiments to observe the corresponding triple transition in solid HD are currently in preparation.

aM. Mengel, B.P. Winnewisser, and M. Winnewisser, J. Mol. Spectrosc., in press, (1998).

bM. Mengel, B.P. Winnewisser, and M. Winnewisser, J. Low Temp. Phys. (Proceedings of the Cryocrystal ’97 Conference), in press, (1998).

cR.H. Tipping, Q. Ma, and J.D. Poll, Phys. Rev. B 22, 12314–12317 (1991).